A small startup company produces lens and apertures for DSLR camera that they sell through a website. After extensive research, the company has developed a revenue function as follows

R(x,y) = x(110-4.5x) + y(155-2y)

in thousand ringgit where x is the number of lens produced and sold in thousands and y is the number of aperture produced and sold in thousands. The corresponding cost function is

C (x,y) = 3x2 + 3y2 + 5xy – 5y + 50

in thousand ringgit. If you are the investors, find the production levels that maximize the revenue. Then, would you invest or not in the company? Justify
1
Expert's answer
2019-12-12T12:39:41-0500

"R(x,y)=x(110-4.5x) + y(155-2y)"


Now we differentiate the function partially with respect to x and y and equate it to 0, in order to find the production levels for maximum revenue generation.


"\implies \partial R/\partial x =110-9x=0"

"\implies x=110/9"

Similarly, "\partial R/\partial y=155-4y=0"

"\implies y=155/4"


Calculating revenue for these production levels, we get the maximum possible revenue, which is;

"R(110\/9,155\/4)=3675.347"


The cost function is given by

"C (x,y) = 3x^2 + 3y^2 + 5xy \u2013 5y + 50"

For the same above obtained production levels, the cost is

"C(110\/9,155\/4)=7177.141"


Clearly, the company cannot be invested in because even for production levels generating maximum revenue, the cost of production is more than the revenue generated, thus resulting in a net loss and thus making the investment not feasible.